You are the Manager of a new computer manufacturing company, Fongkomputer. Tests have shown that the mean life time of your computers is normally distributed, and on average your computers last for only 9 months, with a standard deviation of 3 months. The managing director wants the chance of computer being returned within its guarantee period to be 10%. Suggest a guarantee period that will achieve this objective. Round your solution to the nearest month


1
Expert's answer
2022-05-23T12:22:47-0400
"P(X<x)=P(Z<\dfrac{x-9}{3})=0.1""\dfrac{x-9}{3} =-1.28155"

"x=5"

5 months.